OpenStudy (anonymous):

Given the function f(x)=6x^2+7x-8 Calculate the for value f(x+1) I must be missing a step because I keep coming up with the wrong answer

hero (hero):

\[6(x+1)^2+7(x+1)-8 =6(x^2 +2x+1) + 7x+7-8 = 6x^2 +12x + 6 + 7x + 7 - 8 = 6x^2 + 19x - 2\]

OpenStudy (zarkon):

almost

OpenStudy (zarkon):

6+7-8

OpenStudy (zarkon):

now the coefficient next to the x is wrong

hero (hero):

That's impossible

OpenStudy (anonymous):

\[f(x)=6x^{2}+7x-8\]\[f(x+1)=6(x+1)^{2}+7(x+1)-8\]\[f(x+1)=6(x^{2}+2x+1)+7x+7-8\]\[f(x+1)=6x^{2}+12x+6+7x+7-8\]\[f(x+1)=6x^{2}+19x+5\]

hero (hero):

That's what I had. Darn it

hero (hero):

Zarkon your interference made it worse, lol

OpenStudy (anonymous):

Hey... I was already writing that when you came. I was only slow and laggy here.

OpenStudy (zarkon):

;)...sorry